A Church Songbook  1  30 May’96

1 — What A Mighty God We Serve [C]

\begin{matrix}c \\ ◯ \end{matrix}Public Domain W&M: Unknown Isaiah 9:6

Hosanna! Music Book I, #93.

{C♯ \atop What} a mighty God we serve, What a mighty God we {G7 \atop serve}, {C♭ \atop An}gels bow before Him, {C \atop Hea}ven and earth adore Him, {C \atop What} a mighty {G7 \atop God} we {C \atop serve!}[F C]

1. O {C \atop Zion,} O {F \atop Zion,} that {G7 \atop bring}est good {C \atop tid}ings, Get thee {F \atop up} into the {G7 \atop High} Moun{C \atop tains} Je{C \atop ru}salem, Je{F \atop ru}salem, that {G7 \atop bring}est good {C \atop tid}ings Lift up thy {F \atop voice} with {G7 \atop all} thy C strength Lift it {F \atop up,} be not afraid; Lift it {C \atop up,} be not afraid Say {Am \atop unto} the {C \atop ci}ties of {G7 \atop Judah,} “Behold your {C \atop God,}C7 Behold your {F \atop God,} Be{C \atop hold} {G7 \atop your} {C \atop God!”}

A Church Songbook  2  30 May’96

1 — What A Mighty God We Serve [D]

{D \atop What} a mighty God we serve, What a mighty God we {A7 \atop serve}, {D \atop An}gels bow before Him, {D \atop Hea}ven and earth adore Him, {D \atop What} a mighty {A7 \atop God} we {D \atop serve!}[G D]

1. O {D \atop Zion,} O {G \atop Zion,} that {A7 \atop bring}est good {D \atop tid}ings, Get thee {G \atop up} to into the {A7 \atop High} Moun{D \atop tains} Je{D \atop ru}salem, Je{G \atop ru}salem, that {A7 \atop bring}est good {D \atop tid}ings Lift up thy {G \atop voice} with {A7 \atop all} thy D strength Lift it {G \atop up} be not afraid, Lift it {D \atop up} be not afraid Say {Bm \atop unto} the {D \atop ci}ties of {A7 \atop Judah,} “Behold your {D \atop God,}D7 Behold your {G \atop God,} Be{D \atop hold} {A7 \atop your} {D \atop God!”}

A Church Songbook  3  30 May’96

2 — Finally, My Brethren [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Ephesians 6:10

, .

1. {D \atop Fi}nally, my brethren, be {C \atop strong} in the {D \atop Lord,} {D \atop Fi}nally, my brethren, be {C \atop strong} in the {D \atop Lord,} And in the {G \atop pow}er {A \atop of} His {D \atop might,} And in the {G \atop pow}er {A \atop of} His {D \atop might.} For it is {G \atop God} at {A \atop work} with{D \atop in} you, Both to {G \atop will} {A \atop and} to {D \atop do,} {G \atop All} his {A \atop glo}rious {D \atop plea}sure. His {G \atop truth} shall {A \atop sur}ely en{D \atop dure.}

2…. strength … 3…. mercy … 4…. grace …

5. {D \atop Fi}nally my brethren be {C \atop strong} in the {D \atop Lord,} {D \atop Fi}nally my brethren be {C \atop strong} in the {D \atop Lord,} And in the {G \atop pow}er {A \atop of} His {D \atop might,} And in the {G \atop pow}er {A \atop of} His {D \atop might,} And in the {G \atop pow}er {A \atop of} His {D \atop might.}

Property of a Church  Last Revised: February 11, 1993  Material used by permission.

A Church Songbook  4  30 May’96

3 — The Lord Of Covenant [Am]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

{Am \atop Lift} your hands, for {E \atop He} is Holy. Dm7 Worship Him in {E \atop spir}it and {Am \atop truth.} For His face, is {E \atop like} the sun, {Dm7 \atop The} Lord of {E \atop Cove}{Am \atop nant.}

1. {Dm \atop Clap} your hands and {G \atop shout} in victo{C \atop ry,} For the Lamb was {E \atop slain} but lives. {Dm7 \atop Dance} for joy and {G \atop give} Him thanks. (Men) For the {C \atop free}dom … (Women) For the freedom His {E \atop life} gives. (Men) For the {F \atop free}dom … (Women) For the freedom His {E \atop life} gives.

A Church Songbook  5  30 May’96

2. {Dm \atop Come} you people, {G \atop lift} your voices, C Praise Him with your {E \atop hearts} as one. {Dm7 \atop There} is One who {G \atop turn}ed the key. (Men) And the {C \atop door} … (Women) To His throne is {E \atop open.} (Men) Yes, the {F \atop door} … (Women) To His throne is {E \atop open.}

3. {Dm \atop He} creates Je{G \atop rusa}lem for re{C \atop joic}ing, For re{E \atop joic}ing, {Dm7 \atop And} her people {G \atop for} gladness.

(Men) {C \atop He} …

(Women) He is glad in His {E \atop peo}ple. (Men) Yes, {F \atop He} … (Women) He is glad in His {E \atop peo}ple.

A Church Songbook  6  30 May’96

4 — We Will Rejoice [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Song Of Solomon 1:4

Hosanna! Music Book II, #143.

{D \atop We} will rejoice in {G \atop You} and be glad, {D \atop We} will extol Your {A7 \atop love} more than wine {D \atop Draw} me after {D7 \atop You} and let us {G \atop run} together, {D \atop We} will rejoice in {G \atop You} and {A7 \atop be} {D \atop glad}

Ch: {D \atop Lift} up the light of Thy {G \atop coun}tenance, U{D \atop pon} us O {A7 \atop Lord} {D \atop Lift} up the light of Thy {G \atop coun}tenance, U{D \atop pon} {A7 \atop us} O {D \atop Lord} G D

A Church Songbook  7  30 May’96

5 — Praise Him In His Sanctuary [Em]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Psalm 150

, .

1. Praise Him {Em \atop in} His sanctuary Praise Him {Em \atop in} the skies above Praise Him {G \atop for} the acts of {D \atop pow}er that He {Em \atop does} Praise Him {Em \atop for} surpassing greatness With the {Em \atop trum}pet, harp and lyre With the {G \atop tam}bourines and {Am \atop dan}cing, praise Him {Em \atop now}

Come and {G \atop praise} {D \atop Him} {Bm \atop for} the Lord is {Em \atop good} And His mercy is {B7 \atop ever}Em lasting Come and {G \atop praise} {D \atop Him} {Bm \atop for} the Lord is {Em \atop good} And His mercy is {B7 \atop ever}Em lasting

2. Praise Him {Em \atop with} the clashing cymbals Let them {Em \atop hear} it far and near With the G strings and flutes we’ll {D \atop praise} the Lord our {Em \atop God} Who with {Em \atop majes}ty is reigning He has {Em \atop power} over all Give Him {G \atop glory} and be Am thankful for His {Em \atop love}

Property of a Church  Last Revised: February 11, 1993  Material used by permission.

A Church Songbook  8  30 May’96

6 — Rise Up [Em]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

Rise {Em \atop up,} rise up, we are the Am soldiers of the {Em \atop cross} We are the {Am \atop ones} who are to glorify the {B7 \atop King} Cre{Em \atop ation} groans for the {Am \atop Sons} of God to {Em \atop come} Mani{Am \atop fest}ing all the glory of the {B7 \atop King} So {Am \atop rise} and shine, for your {Em \atop light} has come, And the {Am \atop glor}y of the Lord is u{Em \atop pon} you Lift {Am \atop up} your eyes round a{Em \atop bout} and see All the Am nations are falling at your {B7 \atop feet}

Ch: So let’s {Em \atop glor}ify the Lord, Let’s {Am \atop glor}ify the {Em \atop King} Let’s shout and sing as we {Am \atop take} the victo{B7 \atop ry} So let’s {Em \atop glor}ify the Lord, Let’s {Am \atop glor}ify the {Em \atop King} Let’s {Am \atop shout} and {B7 \atop take} the victo{Em \atop ry!}

A Church Songbook  9  30 May’96

7 — We Are Marching In Messiah’s Band [Am]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

1. We are {Am \atop march}ing, in Messiah’s band; The keys of Am vict’ry in His mighty hand; Let us {Dm \atop go} on to take the promised {Am \atop land.} Raise the Am anthem, sing the victory song; Praise the {Am \atop Lord} for the battle’s won; No {Dm \atop wea}pon formed against Him shall {Am \atop stand.}

A Church Songbook  10  30 May’96

For the {F \atop Cap}tain of the Host is {C \atop Je}sus; We are {F \atop fol}lowing in His {C \atop foot}steps. No {F \atop foe} can stand aAm gainst us in the {F \atop fray.} G 8 — A Song Of Love [G]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

A {G \atop Song} of {Em \atop Love,} I {Am7 \atop sing} to {D \atop You}

A {G \atop Song} of {Em \atop ado}ration, {Am7 \atop to} the {D \atop King}

You {Am7 \atop are} the {D \atop Lord,} the {G \atop Ma}ker of^{G/F♯} the Em universe

You {Am7 \atop are} the {D \atop King} of all {G \atop Kings}

A Church Songbook  11  30 May’96

9 — Hail To The King [D]

\begin{matrix}c \\ ◯ \end{matrix}1989 Christopher Rath W&M: Christopher Rath

New Wine Covenant Church Song Book, #70.

1. I {D \atop praise} You, Lord, for {C \atop You} are {G \atop my} {D \atop King} I {D \atop praise} You, Lord, for {C \atop You} are {G \atop my} {D \atop King}

All {C \atop hail} toG the {D \atop King,} All {C \atop hail} toG King Je{D \atop —}sus All {C \atop hail} toG the {D \atop King} of Kings, All {C \atop hail} toG the {D \atop King}

2. We {D \atop praise} You, Lord, For {C \atop You} are {G \atop our} {D \atop King} We {D \atop praise} You, Lord, For {C \atop You} are {G \atop our} {D \atop King}

3. I {D \atop wor}ship You, my {C \atop Lord} and {G \atop my} {D \atop God} I {D \atop wor}ship You, my {C \atop Lord} and {G \atop my} {D \atop God}

Property of a Church  Last Revised: November 22, 1993  Material used by permission.

A Church Songbook  12  30 May’96

10 — The Horse And Rider [C]

\begin{matrix}c \\ ◯ \end{matrix}1950 Mills Music (music only) W&M: Anonymous and I. Miron & J. Grossman Exodus 15:1–2

Songs Of Praise, #214.

1. {C \atop I} will sing unto the Lord, for {F \atop He} has triumphed Dm gloriously The {G7 \atop horse} and rider thrown into the {C \atop sea} {C \atop I} will sing unto the Lord, for {F \atop He} has triumphed Dm gloriously The {G7 \atop horse} and rider thrown into the {C \atop sea}

2. The {C \atop Lord,} my God, my F strength, my {Dm \atop song,} Is {G7 \atop now} become my victo{C \atop ry} The {C \atop Lord,} my God, my F strength, my {Dm \atop song,} Is {G7 \atop now} become my victo{C \atop ry}

3. The {C \atop Lord} is God and {F \atop I} will Dm praise Him, My G7 Father’s God, and {C \atop I} will ex{G7 \atop alt} Him! The {C \atop Lord} is God and {F \atop I} will Dm praise Him, My G7 Father’s God, and {C \atop I} will ex{C \atop alt} Him!

11 — Nothing But The Blood [D]

\begin{matrix}c \\ ◯ \end{matrix}Public Domain W&M: Robert Lowry Ephesians 1:7

Hosanna! Music Book I, #65.

1. {D \atop What} can wash a{F♯m \atop way} my {Bm \atop sin?}

{G \atop Noth}ing but the {D \atop blood} of {A7 \atop Je–}{D \atop sus.}

{D \atop What} can make me {F♯m \atop whole} a{Bm \atop gain?}

{G \atop Noth}ing but the {D \atop blood} of {A7 \atop Je–}{D \atop sus.}

{D \atop Oh,} Precious {F♯m \atop is} the {Bm \atop flow}

{G \atop That} {Em \atop makes} me {A \atop white} as {D \atop snow.}A7

{D \atop No} other {F♯m \atop fount} I {Bm \atop know,}

{G \atop Noth}ing but the {D \atop blood} of {A7 \atop Je–}{D \atop sus.}

2. {D \atop This} is all my F♯m ; righteous{Bm \atop ness,}

{G \atop Noth}ing but the {D \atop blood} of {A7 \atop Je–}{D \atop sus.}

{D \atop This} is all my {F♯m \atop hope} and {Bm \atop peace,}

{G \atop Noth}ing but the {D \atop blood} of {A7 \atop Je–}{D \atop sus.}

A Church Songbook  14  30 May’96

12 — Far Above All Rule And Authority [C]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Ephesians 1:20–21; 2:6

, .

Far a{C \atop bove} all rule and au{F \atop thor}ity, and {C \atop po}wer and do{G7 \atop min}{C \atop ion} Far a{C \atop bove} all rule and au{F \atop thor}ity, and {C \atop po}wer and do{G7 \atop min}{C \atop ion} And every {F \atop name} that is {G7 \atop nam}ed, not {C \atop on}ly in this age But {F \atop al}so in the {G7 \atop one} to {C \atop come} And every {F \atop name} that is {G7 \atop nam}ed, not {C \atop on}ly in this age But {F \atop al}so in the {G7 \atop one} to {C \atop come} Christ is {C \atop seat}{G \atop ed} at the {C \atop Fa}ther’s right hand, In {F \atop hea}{G7 \atop —–}venly {C \atop pla}ces. We are {C \atop seat}ed with {G \atop Him,} at the {C \atop Fa}ther’s right hand, In {F \atop hea}{G7 \atop —–}venly {C \atop pla}ces.

A Church Songbook  15  30 May’96

13 — Sing Hallelujah [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Scripture In Song Volume II, #167.

{D \atop Sing} Halle{G \atop lu}jah, {D \atop sing} halle{A7 \atop lujah,} {D \atop sing} halle{G \atop lu}jah, there’s {D \atop joy} {A7 \atop in} the {D \atop Lord}

1. There {D \atop is} a king in {G \atop Zion,} There {D \atop is} a king in {A7 \atop Zion,} There {D \atop is} a king in {G \atop Zion}, The {D \atop Ci}ty {A7 \atop of} our {D \atop God!}

2. There’s {D \atop oil} and wine in {G \atop Zion,} There’s {D \atop oil} and wine in {A7 \atop Zion,} There’s {D \atop oil} and wine in {G \atop Zion,} The {D \atop Ci}ty {A7 \atop of} our {D \atop God!}

3. {D \atop Let’s} go up to {G \atop Zion,} {D \atop Let’s} go up to {A7 \atop Zion,} {D \atop Let’s} go up to {G \atop Zion}, The {D \atop Ci}ty {A7 \atop of} our {D \atop God!}

A Church Songbook  16  30 May’96

4. The {D \atop church} is built on {G \atop Zion,} The {D \atop church} is built on {A7 \atop Zion} The {D \atop church} is built on {G \atop Zion,} The {D \atop City} {A7 \atop of} our {D \atop God!}

{D \atop Sing} Halle{G \atop lu}jah, {D \atop sing} halle{A7 \atop lujah,} {D \atop sing} halle{G \atop lu}jah, there’s {D \atop joy} {A7 \atop in} the {D \atop Lord}

A Church Songbook  17  30 May’96

14 — How Lovely Is Your Dwelling Place [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

How {D \atop love}ly is Your {G \atop dwell}ing place, D O Lord of {G \atop Hosts.} My {D \atop soul} longs and {G \atop yearns} for Your {A7 \atop courts,} And my {D \atop heart} and {G \atop flesh} sing for {A7 \atop joy} to the living {D \atop God.} G D One {D \atop day} in Your {G \atop pres}ence is far {D \atop bet}ter to me than {G \atop gold} Or to {D \atop live} my {G \atop whole} life somewhere {Em \atop else.} A7 And {D \atop I} would rather {G \atop be} a {D \atop door} keeper in Your {G \atop house} Than to {D \atop take} my {G \atop fate} upon my{Em \atop self.} A7 You are my {D \atop Sun} and my {G \atop Shield,} You’re my {D \atop Lov}er from the {G \atop start,} And the {D \atop high}way to Your {A7 \atop city} runs through my {G \atop heart.} D

A Church Songbook  18  30 May’96

15 — Zephaniah 3:17 [G]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

G The Lord your {D \atop God} is {Em \atop in} your midst, G The Lord of {D \atop lords,} who sa{Em \atop ves,} G He will ex{D \atop ult} over {Em \atop you} with joy, C He will renew you {D \atop in} His love, C He will rejoice over {D \atop you,} With shouts of {G \atop joy,} D Em with shouts of {G \atop joy,} D Em With shouts of {C \atop joy,} D with shouts of {C \atop joy,} D With shouts of {G \atop joy!} 16 — Come And Let Us Go [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Micah 4:2

, .

{D \atop Come} and let us {Bm \atop go} to the {Em \atop moun}tain of the {A7 \atop Lord,}

And {Em \atop to} the {A7 \atop house} of our {D \atop God.}

{D \atop Come} and let us {Bm \atop go} to the {Em \atop moun}tain of the {A7 \atop Lord,}

And {Em \atop to} the {A7 \atop house} of our {D \atop God.}

And {G \atop He} will {A \atop teach} us of His {D \atop ways,} D7

And {G \atop we} will {A \atop walk} in His {D \atop paths.} D7

And the {G \atop law} shall go {A \atop forth} from {D \atop Zi}{Bm \atop on,}

And the {Em \atop Word} of the {A \atop Lord} from Jerusa{D \atop lem.} 17 — Behold, I Am The Lord [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Sounds Of Zion, Volume III.

Be{D \atop hold,} I am the {F♯m \atop Lord,} the {G \atop God} of all {A \atop flesh.}

Is there {D \atop any}thing,D7 is there {G \atop any}thing,Em {Gm \atop too} {D \atop hard} {A \atop for} {D \atop Me?}

Is there {D \atop any}thing, {G \atop any}thing, {D \atop any}thing, too hard {G \atop for} {A \atop Me?}

{A7 \atop Is} there {D \atop any}thing, {G \atop any}thing, {D \atop any}thing,Bm7 {Gm \atop too} hard^{D/A} {A \atop for} {D \atop Me?}

A Church Songbook  21  30 May’96

18 — Stand Up And Bless [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Nehemiah 9:5

Hosanna! Music Book IV, #332.

Stand {D \atop up} and bless the {Em \atop Lord} your God From ever{A \atop last}ing to ever{D \atop last}ing Stand {D \atop up} and bless the {Em \atop Lord} your God From ever{A \atop last}ing to ever{D \atop last}ing

Ch: And {G \atop ble}ssed be your glorious {D \atop name,} O Lord Which is eEm xaltedA above all {D \atop ble}ssing and {D7 \atop praise} And {G \atop ble}ssed be your glorious {D \atop name,} O Lord Which is eEm xalted,A which is exal—^{G/D}{D \atop ted}

A Church Songbook  22  30 May’96

19 — Jesus Is The Lord [G]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

{G \atop Je}sus is the {G7 \atop Lord} {C \atop Je}sus the Lord Am reigns {D \atop We} shall take the {D7 \atop king}doms of this {G \atop world} {C \atop in} His {G \atop name} D7 {G \atop Ev}ery tribe and {G7 \atop nation} {C \atop Ev}ery situ{Am \atop ation} {D \atop Must} declare that {D7 \atop Jesus} is the {G \atop Lord} G7

Ch: For the {C \atop Lord} {G \atop our} {D \atop God} has de{G \atop liver}ed Him from {Em \atop death} And esAm tablished {D \atop Je}sus as {G \atop Lord} G7 He has {C \atop gi}ven {G \atop Him} the {D \atop po}wer over {G \atop all} that He has {Em \atop made} For our {Am \atop God} has {D \atop made} His Christ the {G \atop Lord}

A Church Songbook  23  30 May’96

20 — I Will Sing Praises [G]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

G I will sing {C \atop prais}es, G I will sing {C \atop prais}es, G I will sing {C \atop prais}es {D \atop to} my {G \atop God} G I will sing {C \atop prais}es, G I will sing {C \atop prais}es, G I will sing {C \atop prais}es {D \atop to} my {G \atop God} G And when He {C \atop comes} a{D \atop gain,} G The rocks will not {C \atop have} to {D \atop cry} out G His people will {C \atop shout} the mighty {D \atop name} Of the King of {Em \atop Glor}y G When the mountains {C \atop sing} forth G And all the trees {C \atop clap} their hands G His people will {C \atop shout} the name Of the {D \atop beau}ty of holiness: {D7 \atop Jesus} Christ is His {G \atop name!}

A Church Songbook  24  30 May’96

21 — Let God Be Magnified [G]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

Let {G \atop all} those that seek Thee rejoice and be glad In {Am \atop Thee}, In {D7 \atop Thee} And {G \atop let} such as love thy salvation, say Continual{Am \atop ly,} continual{D7 \atop ly,} C“Let God be magnified, Glet God be magnified, {Am \atop Let} God be magni{D7 \atop fied.”} C“Let God be magnified, Glet God be magnified, {Am \atop Let} God be {D7 \atop mag}ni{G \atop fied.”}C G

A Church Songbook  25  30 May’96

22 — There’s A Light Shining Forth [D]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

There’s a light shining {D \atop forth,} I can see it on the ho{Em \atop rizon;} A7 It’s the army of {Em \atop God,} A preparing for {D \atop war;} Coming conquering vicF♯m ; torious, D7 O’er the army of {G \atop sa}tan; Em Nothing shall {D \atop stand,}A7 before the army of {D \atop God.} G D

A Church Songbook  26  30 May’96

23 — How Great Is Our God [D]

\begin{matrix}c \\ ◯ \end{matrix}Public Domain W&M: Unknown

Worship Him II, #65.

{D \atop How} great is our {A7 \atop God!} How great is His {D \atop name}G D How great is our {A7 \atop God!} Forever the {D \atop same}G D He rolled back the {D \atop wa}te{D7 \atop rs,} of the mighty Red {G \atop Sea} And He said, “I’ll never {D \atop leave} you,A7 Put your trust in {D \atop Me!”} [G D]

23 — How Great Is Our God [E]

{E \atop How} great is our {B7 \atop God!} How great is His {E \atop name}A E How great is our {B7 \atop God!} Forever the {E \atop same}A E He rolled back the {E \atop wa}te{E7 \atop rs,} of the mighty Red {A \atop Sea} And He said, “I’ll never {E \atop leave} you,B7 Put your trust in {E \atop Me!”} [A E]

A Church Songbook  27  30 May’96

24 — Lift Jesus Higher [E]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

, .

Lift Jesus {E \atop high}{A \atop er.} E Lift Jesus {E \atop high}{A \atop er.} E Lift Him up for the world to {B7 \atop see.} He said, “If {E \atop I} E7 be lifted {A \atop up} from the earth, I will {E \atop draw} all {B7 \atop men} unto {E \atop me}.”

25 — Garment Of Praise [C]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Isaiah 61:3

, .

I have {C \atop put} on my garment of {G7 \atop praise,}

I have {G7 \atop put} on my garment of {C \atop prai}{C7 \atop se}

And the {F \atop spirit} of heaviness, is {C \atop gone} from {Am \atop me.}

I have {C \atop put} on my {G7 \atop gar}ment of {C \atop praise!} [F C]

25 — Garment Of Praise [D]

I have {D \atop put} on my garment of {A7 \atop praise,}

I have {A7 \atop put} on my garment of {D \atop prai}{D7 \atop se}

And the {G \atop spirit} of heaviness, is {D \atop gone} from {Bm \atop me.}

I have {D \atop put} on my {A7 \atop gar}ment of {D \atop praise!} [G D] 26 — I Was Glad [C]

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Psalm 122:1

, .

I was {C \atop glad,} very glad, when they {F \atop said} unto me

Let us {G7 \atop go,} into the house of the {C \atop Lord,} today

There is singing, there is {C7 \atop danc}ing, there is {F \atop vic}tory

In the {G7 \atop house} of the Lord, to{C \atop day!} 27 — Oh, The Blood [C]

\begin{matrix}c \\ ◯ \end{matrix}Public Domain W&M: Unknown Psalm 51:7

Hosanna! Music Book I, #68.

{C \atop Oh} the blood {F \atop of} {C \atop Jesus.}

{G7 \atop Oh} the blood of {C \atop Jesus.}

Oh the blood {F \atop of} {C \atop Je–}{Am \atop sus.}

It {Dm \atop wash}es {G7 \atop white} as {C \atop snow.}

27 — Oh, The Blood [D]

{D \atop Oh} the blood {G \atop of} {D \atop Jesus.}

{A7 \atop Oh} the blood of {D \atop Jesus.}

Oh the blood {G \atop of} {D \atop Je–}{Bm \atop sus.}

It {Em \atop wash}es {A7 \atop white} as {D \atop snow.}

27 — Oh, The Blood [E]

{E \atop Oh} the blood {A \atop of} {E \atop Jesus.}

{B7 \atop Oh} the blood of {E \atop Jesus.}

Oh the blood {A \atop of} {E \atop Je–}{C♯m \atop sus.}

It {F♯m \atop wash}es {B7 \atop white} as {E \atop snow.}

27 — Oh, The Blood [F]

{E♭ \atop Oh} the blood {A♭ \atop of} {E♭ \atop Jesus.}

{B♭7 \atop Oh} the blood of {E♭ \atop Jesus.}

Oh the blood {A♭ \atop of} {E♭ \atop Je–}{Cm \atop sus.}

It {Fm \atop wash}es {B♭7 \atop white} as {E♭ \atop snow.}